trace the conic x2+2xy+y2-3x+y-2=-0


1
Expert's answer
2021-01-14T17:57:35-0500

General Cartesian form of the equation is

"Ax^2\n +Bxy+Cy^2\n +Dx+Ey+F=0;"

If discriminant is equal to zero, the equation represents a parabola.

"B^2 - 4AC = 2^2 - 4\times1\times1 = 0."

Therefore it is a parabola.


"x^2+2xy+y^2-3x+y-2=0"


is a parabola with vertex at "(h,k)=(-\frac{3}{4},-\frac{1}{2})"

and focal length  "\left|p\right|=\frac{3}{4}"

"4\cdot (\frac{3}{4})(y-(-\frac{3}{4}))=(x-(-\frac{1}{2}))^2" is the standand equation of the parabola

Focus : "(0,-\frac{1}{2})"

directrix: "x = - \frac{3}{2}"

vertex: "(-\frac{3}{4},-\frac{1}{2})"

Axis: horizontal

Standard Form:

"4\cdot (\frac{3}{4})(y-(-\frac{3}{4}))=(x-(-\frac{1}{2}))^2"

Below is the graph of conic